# A system of NLS arising in optical material without Galilean symmetry

### v1

@@ 2 Existence and non-existence of traveling solitary waves
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If a sequence $\{(u_{n},v_{n})\}$ in $\tilde{X}_{\omega,c}$ satisfies
[…]
for some positive constants $A_{1},A_{2}>0$, then there exist $\{y_{n}\}$ and $(u,v)\in\tilde{X}_{\omega,c}\setminus\{(0,0)\}$ such that $\{\tilde{\tau}_{y}(u_{n},v_{n})\}$ has a subsequence that weakly converges to $(u,v)$ in $\tilde{X}_{\omega,c}$.
@@
Since $\tilde{Q}_{\omega,c}(u_{n},v_{n})\to A_{1}$, we deduce that the sequence $\{(u_{n},v_{n})\}$ is bounded in $\tilde{X}_{\omega,c}$. Since $\tilde{D}(u_{n},v_{n})\to A_{2}$, we obtain
@@
Hence, by [13, Lemma 6], we can obtain the desired result. ∎
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If a sequence $\{(u_{n},v_{n})\}$ in $\tilde{X}_{\omega,c}$ satisfies
[…]
then there exist $\{y_{n}\}$ and $(u,v)\in\tilde{X}_{\omega,c}\setminus\{(0,0)\}$ such that $\{\tilde{\tau}_{y_{n}}(u_{n},v_{n})\}$ has a subsequence that converges to $(u,v)$ in $\tilde{X}_{\omega,c}$. In particular, $(u,v)\in\tilde{a}_{\omega,c}$.
@@
$\frac{1}{2}\tilde{Q}_{\omega,c}(u_{n},v_{n})=\tilde{S}_{\omega,c}(u_{n},v_{n})-\frac{1}{4}\tilde{N}(u_{n},v_{n})\to\tilde{\mu}_{\omega,c},$
$\tilde{D}(u_{n},v_{n})=\tilde{S}_{\omega,c}(u_{n},v_{n})-\frac{1}{2}\tilde{N}_{\omega,c}(u_{n},v_{n})\to\tilde{\mu}_{\omega,c}.$
By Lemma 2.6 and Lemma 2.9, there exist $\{y_{n}\}$, $(u,v)\in\tilde{X}_{\omega,c}\setminus\{(0,0)\}$, and a subsequence of $\tilde{\tau}_{y_{n}}(u_{n},v_{n})$(still denoted by $\tilde{\tau}_{y_{n}}(u_{n},v_{n})$) such that $\tilde{\tau}_{y_{n}}(u_{n},v_{n})\rightharpoonup(u,v)$ weakly in $\tilde{X}_{\omega,c}$.
@@
$\tilde{Q}_{\omega,c}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))-\tilde{Q}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to\tilde{Q}_{\omega,c}(u,v),$ (2.9)
$\tilde{N}_{\omega,c}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))-\tilde{N}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to\tilde{N}_{\omega,c}(u,v).$ (2.10)
From (2.9) and $\tilde{Q}_{\omega,c}(u,v)>0$, we obtain that, up to subsequence,
$\frac{1}{2}\lim_{n\to\infty}\tilde{Q}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)<\frac{1}{2}\lim_{n\to\infty}\tilde{Q}_{\omega,c}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))=\tilde{\mu}_{\omega,c}.$
From this and Lemma 2.7, we obtain $\tilde{N}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)>0$ for $n$ large enough. Therefore, from (2.10), $\tilde{N}_{\omega,c}(u,v)\leq 0$ since $\tilde{N}_{\omega,c}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))\to 0$. Again, by Lemma 2.7 and the weakly semi-continuity of norms, we have
$\tilde{\mu}_{\omega,c}\leq\frac{1}{2}\tilde{Q}_{\omega,c}(u,v)\leq\frac{1}{2}\tilde{Q}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})\right)=\tilde{\mu}_{\omega,c}.$
Therefore, by (2.9), we deduce $\tilde{Q}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to 0$ as $n\to\infty$, which implies that $\tilde{\tau}_{y_{n}}(u_{n},v_{n})\to(u,v)$ strongly in $\tilde{X}_{\omega,c}$. Now we have completed the proof of this lemma. ∎
@@
Combining the Lemmas 2.4, 2.5 and 2.10, one can get Proposition 2.2. ∎
In order to prove Theorem 1.1 (ii), we will prove that the minimizer $(u,v)$ is a non-semitrivial solution ($u\neq 0$ and $v\neq 0$) of the system (1.9). If $u=0$, one can take $v=\tilde{Q}$, where $\tilde{Q}$ is a Galilean transformation of $Q$, and $Q$ is the positive ground state of equation (1.13), then $(0,\tilde{Q})$ is the semi-trivial solution of system (1.9). Since $Q$ is the ground state of (1.13), then $S_{\omega,0}(Q)\leq S_{\omega,c}(\tilde{Q})$. As below, we give the relation between $S_{\omega,c}(u,v)$ and $S_{\omega,0}(0,Q)$.
@@
Assume that $3\gamma>\left(\frac{9}{\sqrt{2}}\right)^{\frac{4}{N-4}}$. Then the functions $(u,v)$ in the Nehari manifold $N_{\omega,c}$ satisfy
$S_{\omega,c}(u,v)<S_{\omega,c}(0,Q)=S_{\omega,0}(0,Q),$
where $Q$ is the positive ground state of (1.13). In particular, $(0,Q)$ is not a minimizer of $S_{\omega,c}$ in the complex space and $(u,v)$ is a non-semitrivial solution ($u\neq 0$ and $v\neq 0$).
@@
It suffices to prove that the existence of $\alpha_{1},\alpha_{2}\in\mathbb{R}$ and the complex function $h\in H^{1}$ such that $(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)\in\mathcal{N}_{\omega,c}$ and $S_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)<S_{\omega,c}(0,Q)$. In fact, $(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)\in\mathcal{N}_{\omega,c}$ if and only if $N_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=0$. Since
$N_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=2Q_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)-4D(\alpha_{1}\alpha_{2}h,\alpha_{1}Q),$
by taking $\alpha_{1}\in\mathbb{R}$ satisfying
$\alpha_{1}^{2}=\frac{Q_{\omega,c}(\alpha_{2}h,Q)}{2D(\alpha_{2}h,Q)}.$ (2.11)
We see that $N_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=0$. From this point, we take $\alpha_{1}$ as in (2.11).
@@
$Q_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=2D(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)$
@@
$S_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=$ $Q_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)-D(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)$
$=$ $\frac{1}{2}Q_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)$
$=$ $\frac{\alpha_{1}^{2}}{2}Q_{\omega,c}(\alpha_{2}h,Q)$
$=$ $\frac{Q_{\omega,c}^{2}(\alpha_{2}h,Q)}{4D(\alpha_{2}h,Q)}.$
@@
$S_{\omega,c}(0,Q)=$ $\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega}{2}\|Q\|_{L^{2}}^{2}+\frac{1}{2}c\cdot P(0,Q)-\frac{9}{4}\int|Q|^{4}$
$=$ $\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega}{2}\|Q\|_{L^{2}}^{2}-\frac{9}{4}\int|Q|^{4}$
$=$ $\frac{1}{4}Q_{\omega,0}(0,Q),$
since $Q$ is a real function and $P(0,Q)=0$. Thus $S_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)<S_{\omega,0}(0,Q)$ is equivalent to
$Q_{\omega,c}^{2}(\alpha_{2}h,Q)<4D(\alpha_{2}h,Q)Q_{\omega,0}(0,Q).$ (2.12)
Both sides of (2.12) are polynomials of degree $4$ in $\alpha_{2}$. The leading coefficient of the polynomial in the left-hand side is
$\left(\frac{1}{2}\int|\nabla h|^{2}+\frac{\omega}{2}|h|^{2}\right)^{2}$
@@
$\frac{1}{9}\int|h|^{4}\left(\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega}{2}|Q|^{2}\right).$
Therefore, (2.12) holds for $\alpha_{2}$ sufficiently large, provided that
$Q_{\omega,c}^{2}(h,0)<\frac{1}{9}\|h\|_{L^{4}}^{4}Q_{\omega,0}(0,Q).$ (2.13)
@@
$\left(\frac{\lambda^{2}}{2}\int|\nabla Q|^{2}+\frac{\omega}{2}\int|Q|^{2}\right)^{2}<\frac{\lambda^{N}}{9}Q_{\omega,0}(0,Q)\|Q\|_{L^{4}}^{4}.$ (2.14)
@@
$\int|\nabla Q|^{2}=\frac{3\gamma\omega N}{4-N}\int|Q|^{2}=\frac{9N}{4}\int|Q|^{4}.$
@@
$\left(\frac{\lambda^{2}}{2}\frac{3\gamma\omega N}{4-N}\int|Q|^{2}+\frac{\omega}{2}\int|Q|^{2}\right)^{2}<\frac{\lambda^{N}}{9}\left(\frac{1}{2}\frac{3\gamma\omega N}{4-N}\int|Q|^{2}+\frac{3\gamma\omega}{2}\int|Q|^{2}\right)\frac{12\gamma\omega}{9(4-N)}\int|Q|^{2}.$ (2.15)
@@
$\left(\frac{\lambda^{2}}{2}\frac{3\gamma\omega N}{4-N}+\frac{\omega}{2}\right)^{2}<\frac{\lambda^{N}}{9}\left(\frac{1}{2}\frac{3\gamma\omega N}{4-N}+\frac{3\gamma\omega}{2}\right)\frac{12\gamma\omega}{9(4-N)}.$
That is
$\left(\lambda^{2}3\gamma N+(4-N)\right)^{2}<\frac{32\gamma^{2}\lambda^{N}}{9}.$
@@
$f(\lambda)=\left(\lambda^{2}3\gamma N+(4-N)\right)-\left(\frac{32\gamma^{2}\lambda^{N}}{9}\right)^{\frac{1}{2}}.$
It is easy to see that $f$ has a global minimum at the point $\lambda_{0}=\left(\frac{9}{\sqrt{2}}\right)^{\frac{2}{N-4}}$. In addition, $f(\lambda_{0})=(4-N)(1-3\gamma\lambda_{0}^{2})$. By assumption, $f(\lambda_{0})<0$, which means that $01-3\gamma\lambda_{0}^{2}<$ 0. If we choose that $3\gamma>\left(\frac{9}{\sqrt{2}}\right)^{\frac{4}{N-4}}$, then we see that (2.15) holds for $\lambda=\lambda_{0}$ and the proof of this Lemma is complete. ∎
@@
$(4-N)\int(|\nabla u_{\omega}|^{2}+|\nabla v_{\omega}|^{2})=N\omega(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2})+(N-2)c\cdot P(u_{\omega},v_{\omega}).$ (2.16)
Case 1: $N=2$. In this case, by the identity (2.16), we have
$\int(|\nabla u_{\omega}|^{2}+|\nabla v_{\omega}|^{2})=\omega(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2}).$
If $\omega\leq 0$, it follows that
@@
Combining with identity (2.16), then
$\left(4-N-\frac{|c|}{2}\right)\int|\nabla u_{\omega}|^{2}+\left(4-N-\frac{\gamma|c|}{6}\right)\int|\nabla v_{\omega}|^{2}\leq\left(wN+\frac{|c|}{2}\right)(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2}).$
Thus, if $wN+\frac{|c|}{2}\leq 0$, $04-N-\frac{|c|}{2}>$ 0 and $04-N-\frac{\gamma|c|}{6}>$ 0, then it must be $(u_{\omega},v_{\omega})=(0,0)$. Thus, Theorem 1.2 (ii) holds. ∎

### v2

@@ 2 Existence and non-existence of traveling waves
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If a sequence $\{(u_{n},v_{n})\}$ in $\tilde{X}_{\omega,c,\alpha-3\gamma}$ satisfies
$\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})\to A_{1},~{}~{}\tilde{D}(u_{n},v_{n})\to A_{2}~{}~{}\text{as}~{}~{}n\to\infty,$
for some positive constants $A_{1},A_{2}>0$, then there exist $\{y_{n}\}$ and $(u,v)\in\tilde{X}_{\omega,c,\alpha-3\gamma}\setminus\{(0,0)\}$ such that $\{\tilde{\tau}_{y}(u_{n},v_{n})\}$ has a subsequence that weakly converges to $(u,v)$ in $\tilde{X}_{\omega,c,\alpha-3\gamma}$.
@@
Since $\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})\to A_{1}$, we deduce that the sequence $\{(u_{n},v_{n})\}$ is bounded in $\tilde{X}_{\omega,c,\alpha-3\gamma}$. Since $\tilde{D}(u_{n},v_{n})\to A_{2}$, we obtain
@@
Hence, by [12, Lemma 6], we can obtain the desired result. ∎
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If a sequence $\{(u_{n},v_{n})\}$ in $\tilde{X}_{\omega,c,\alpha-3\gamma}$ satisfies
$\tilde{N}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})\to 0,~{}~{}\tilde{S}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})\to\tilde{\mu}_{\omega,c,\alpha-3\gamma}~{}~{}\text{as}~{}~{}n\to\infty,$
then there exist $\{y_{n}\}$ and $(u,v)\in\tilde{X}_{\omega,c,\alpha-3\gamma}\setminus\{(0,0)\}$ such that $\{\tilde{\tau}_{y_{n}}(u_{n},v_{n})\}$ has a subsequence that converges to $(u,v)$ in $\tilde{X}_{\omega,c,\alpha-3\gamma}$. In particular, $(u,v)\in\tilde{a}_{\omega,c,\alpha-3\gamma}$.
@@
$\frac{1}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})=\tilde{S}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})-\frac{1}{4}\tilde{N}(u_{n},v_{n})\to\tilde{\mu}_{\omega,c,\alpha-3\gamma},$
$\tilde{D}(u_{n},v_{n})=\tilde{S}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})-\frac{1}{2}\tilde{N}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})\to\tilde{\mu}_{\omega,c,\alpha-3\gamma}.$
From Lemma 2.5 and Lemma 2.8, there exist $\{y_{n}\}$, $(u,v)\in\tilde{X}_{\omega,c,\alpha-3\gamma}\setminus\{(0,0)\}$, and a subsequence of $\tilde{\tau}_{y_{n}}(u_{n},v_{n})$(still denoted by $\tilde{\tau}_{y_{n}}(u_{n},v_{n})$) such that $\tilde{\tau}_{y_{n}}(u_{n},v_{n})\rightharpoonup(u,v)$ weakly in $\tilde{X}_{\omega,c,\alpha-3\gamma}$.
@@
$\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))-\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v),$ (2.9)
$\tilde{N}_{\omega,c,\alpha-3\gamma}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))-\tilde{N}_{\omega,c,\alpha-3\gamma}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to\tilde{N}_{\omega,c,\alpha-3\gamma}(u,v).$ (2.10)
From (2.9) and $\tilde{Q}_{\omega,c,\alpha-3\gamma}(u,v)>0$, we obtain that, up to subsequence,
$\frac{1}{2}\lim_{n\to\infty}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)<\frac{1}{2}\lim_{n\to\infty}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))=\tilde{\mu}_{\omega,c,\alpha-3\gamma}.$
Combining this and Lemma 2.6, we obtain $\tilde{N}_{\omega,c,\alpha-3\gamma}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)>0$ for $n$ large enough. Therefore, from (2.10), $\tilde{N}_{\omega,c,\alpha-3\gamma}(u,v)\leq 0$ since $\tilde{N}_{\omega,c,\alpha-3\gamma}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))\to 0$. Again, by Lemma 2.6 and the weakly lower semi-continuity of norms, we have
$\tilde{\mu}_{\omega,c,\alpha-3\gamma}\leq\frac{1}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v)\leq\frac{1}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})\right)=\tilde{\mu}_{\omega,c,\alpha-3\gamma}.$
Therefore, by (2.9), we deduce $\tilde{Q}_{\omega,c,\alpha-3\gamma}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to 0$ as $n\to\infty$, which implies that $\tilde{\tau}_{y_{n}}(u_{n},v_{n})\to(u,v)$ strongly in $\tilde{X}_{\omega,c,\alpha-3\gamma}$. Now we have completed the proof of this lemma. ∎
@@
Combining the Lemmas 2.3, 2.4 and 2.9, one can get Proposition 2.2. ∎
In order to prove Theorem 1.1 (ii), we will prove that the minimizer $(u,v)$ is a non-semitrivial solution ($u\neq 0$ and $v\neq 0$) of the system (1.8). If $u=0$, then $(0,\tilde{Q})$ is the semi-trivial solution of system (1.8), where $\tilde{Q}$ is a Galilean transformation of $Q$, and $Q$ is the positive ground state of equation (1.12). Since $Q$ is the ground state of (1.12), then $S_{\omega,0,\alpha-3\gamma}(Q)\leq S_{\omega,c,\alpha-3\gamma}(\tilde{Q})$. As below, we give the relation between $S_{\omega,c,\alpha-3\gamma}(u,v)$ and $S_{\omega,0,\alpha-3\gamma}(0,Q)$.
@@
Assume that $3\gamma\omega+\alpha-3\gamma>\omega{9}^{\frac{4}{4-N}}$. Then the functions $(u,v)$ in the Nehari manifold $N_{\omega,c,\alpha-3\gamma}$ satisfy
$S_{\omega,c,\alpha-3\gamma}(u,v)<S_{\omega,c,\alpha-3\gamma}(0,Q)=S_{\omega,0,\alpha-3\gamma}(0,Q),$
where $Q$ is the positive ground state of (1.12). In particular, $(0,Q)$ is not a minimizer of $S_{\omega,c,\alpha-3\gamma}$ in the complex space and $(u,v)$ is a non-semitrivial solution ($u\neq 0$ and $v\neq 0$).
@@
It suffices to prove that there exist $\beta_{1},\beta_{2}\in\mathbb{R}$ and the complex function $h\in H^{1}$ such that $(\beta_{1}\beta_{2}h,\beta_{1}Q)\in\mathcal{N}_{\omega,c,\alpha-3\gamma}$ and $S_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)<S_{\omega,c,\alpha-3\gamma}(0,Q)$. In fact, $(\beta_{1}\beta_{2}h,\beta_{1}Q)\in\mathcal{N}_{\omega,c,\alpha-3\gamma}$ if and only if $N_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)=0$. Since
$N_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)=2\mathcal{Q}_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)-4D(\beta_{1}\beta_{2}h,\beta_{1}Q),$
then by taking $\beta_{1}\in\mathbb{R}$ satisfying
$\beta_{1}^{2}=\frac{\mathcal{Q}_{\omega,c,\alpha-3\gamma}(\beta_{2}h,Q)}{2D(\beta_{2}h,Q)}.$ (2.11)
We see that $N_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)=0$. From this point, we take $\beta_{1}$ as in (2.11).
@@
$\mathcal{Q}_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)=2D(\beta_{1}\beta_{2}h,\beta_{1}Q)$
@@
$S_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)=$ $\mathcal{Q}_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)-D(\beta_{1}\beta_{2}h,\beta_{1}Q)$
$=$ $\frac{1}{2}\mathcal{Q}_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)$
$=$ $\frac{\beta_{1}^{2}}{2}\mathcal{Q}_{\omega,c,\alpha-3\gamma}(\beta_{2}h,Q)$
$=$ $\frac{\mathcal{Q}_{\omega,c,\alpha-3\gamma}^{2}(\beta_{2}h,Q)}{4D(\beta_{2}h,Q)}.$
@@
$S_{\omega,c,\alpha-3\gamma}(0,Q)=$ $\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega+\alpha-3\gamma}{2}\|Q\|_{L^{2}}^{2}+\frac{1}{2}c\cdot P(0,Q)-\frac{9}{4}\int|Q|^{4}$
$=$ $\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega+\alpha-3\gamma}{2}\|Q\|_{L^{2}}^{2}-\frac{9}{4}\int|Q|^{4}$
$=$ ${\frac{1}{2}\mathcal{Q}_{\omega,0,\alpha-3\gamma}(0,Q)},$
where we used the fact that $Q$ is a positive, real and radial symmetry function, $P(0,Q)=0$ and the identity
$\int|\nabla Q|^{2}+(3\gamma\omega-3\gamma+\alpha)\int|Q|^{2}=9\int|Q|^{4}.$
Thus $S_{\omega,c,\alpha-3\gamma}(\beta_{1}\beta_{2}h,\beta_{1}Q)<S_{\omega,0,\alpha-3\gamma}(0,Q)$ is equivalent to
$\mathcal{Q}_{\omega,c,\alpha-3\gamma}^{2}(\beta_{2}h,Q)<{2}D(\beta_{2}h,Q)\mathcal{Q}_{\omega,0,\alpha-3\gamma}(0,Q).$ (2.12)
Both sides of (2.12) are polynomials of degree $4$ in $\beta_{2}$. The leading coefficient of the polynomial in the left-hand side is
$\left(\frac{1}{2}\int|\nabla h|^{2}+\frac{\omega}{2}|h|^{2}+\frac{1}{2}c\cdot P(h,0)\right)^{2}$
@@
${\frac{1}{18}}\int|h|^{4}\left(\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega+\alpha-3\gamma}{2}|Q|^{2}\right).$
Therefore, (2.12) holds for $\beta_{2}$ sufficiently large, provided that
$\mathcal{Q}_{\omega,c,\alpha-3\gamma}^{2}(h,0)<{\frac{1}{18}}\|h\|_{L^{4}}^{4}\mathcal{Q}_{\omega,0,\alpha-3\gamma}(0,Q).$ (2.13)
@@
$\left(\frac{\lambda^{2}}{2}\int|\nabla Q|^{2}+\frac{\omega}{2}\int|Q|^{2}\right)^{2}<\frac{\lambda^{N}}{18}\mathcal{Q}_{\omega,0,\alpha-3\gamma}(0,Q)\|Q\|_{L^{4}}^{4}.$ (2.14)
@@
$\int|\nabla Q|^{2}=\frac{(3\gamma\omega+\alpha-3\gamma)N}{4-N}\int|Q|^{2}=\frac{9N}{4}\int|Q|^{4}.$
@@
$\left(\frac{\lambda^{2}}{2}\frac{(3\gamma\omega+\alpha-3\gamma)N}{4-N}\int|Q|^{2}+\frac{\omega}{2}\int|Q|^{2}\right)^{2}$
$<$ $\frac{\lambda^{N}}{18}\left(\frac{(3\gamma\omega+\alpha-3\gamma)N}{2(4-N)}\int|Q|^{2}+\frac{(3\gamma\omega+\alpha-3\gamma)}{2}\int|Q|^{2}\right)\frac{4(3\gamma\omega+\alpha-3\gamma)}{9(4-N)}\int|Q|^{2}.$ (2.15)
@@
$\left(\lambda^{2}(3\gamma\omega+\alpha-3\gamma)N+(4-N)\omega\right)^{2}<\frac{16\lambda^{N}}{81}(3\gamma\omega+\alpha-3\gamma)^{2}.$
@@
$f(\lambda)=\left(\lambda^{2}(3\gamma\omega+\alpha-3\gamma)N+(4-N)\omega\right)-\frac{4\lambda^{\frac{N}{2}}}{9}(3\gamma\omega+\alpha-3\gamma).$
It is easy to see that $f$ has a global minimum at the point $\lambda_{0}={9}^{\frac{2}{N-4}}$. In addition, $f(\lambda_{0})=(4-N)(\omega-(3\gamma\omega+\alpha-3\gamma)\lambda_{0}^{2})$. By assumption, $f(\lambda_{0})<0$, which means that $\omega-(3\gamma\omega+\alpha-3\gamma)\lambda_{0}^{2}<0$. If we choose that $3\gamma\omega+\alpha-3\gamma>\omega{9}^{\frac{4}{4-N}}$, then we see that (2) holds for $\lambda=\lambda_{0}$ and the proof of this Lemma is complete. ∎
@@
$(4-N)\int(|\nabla u_{\omega}|^{2}+|\nabla v_{\omega}|^{2})=$ $N\omega(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2})$
$+N(\alpha-3\gamma)\|v_{\omega}\|_{L^{2}}^{2}+(N-2)c\cdot P(u_{\omega},v_{\omega}).$ (2.16)
Case 1: $N=2$. In this case, by the identity (2), we have
$\int(|\nabla u_{\omega}|^{2}+|\nabla v_{\omega}|^{2})=\omega(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2})+(\alpha-3\gamma)\|v_{\omega}\|_{L^{2}}^{2}.$
If $\omega\leq\min\left\{0,\frac{\alpha-3\gamma}{3\gamma}\right\}$, it follows that
@@
Combining with identity (2), then
$\left(4-N-\frac{|c|}{2}\right)\int|\nabla u_{\omega}|^{2}+\left(4-N-\frac{\gamma|c|}{6}\right)\int|\nabla v_{\omega}|^{2}$
$\leq$ $\left(wN+\frac{|c|}{2}\right)(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2})+N(\alpha-3\gamma)\|v_{\omega}\|_{L^{2}}^{2}.$
Thus, if $wN+\frac{|c|}{2}\leq 0$ and $(3\gamma\omega+\alpha-3\gamma)N+\frac{|c|}{2}<0$, $04-N-\frac{|c|}{2}>$ 0 and $04-N-\frac{\gamma|c|}{6}>$ 0, then it must be $(u_{\omega},v_{\omega})=(0,0)$. Thus, Theorem 1.2 (ii) holds. ∎